OpenStudy (anonymous):

A 2.0 kg purse is dropped from the top of the leaning tower of pisa and falls 55m before reaching the ground with a speed of 29 m/s. What was the average force of air resistance?

OpenStudy (missmob):

First find the net accelereation using v^2 = v0^2 + 2*a*y or a = v^2/2y = 26^2/(2*(-55)) = -6.145m/s^2 the negeative sing means the accelereation is doown apply Newton's Laws m*g - F = m*a So F = m*(g - a) = 2.4*(9.8 - 6.145) = 8.77N

OpenStudy (anonymous):

Another way to solve it is by Energy balance: \[mgh=\frac{ 1 }{ 2}mv^2+Fh \rightarrow F=mg-\frac{ 1 }{ 2}m \frac{ v^2 }{ h}=2·9.81-29^2/55=4.33N\]

OpenStudy (anonymous):

@missmob, review your calculations